OpenStudy (study_buddy99):

am I right? (Alg 2) ( I award medals and maybe a testimonial :)) 3^-3*3^4*3^2=27^3

OpenStudy (study_buddy99):

yay, thank you!

OpenStudy (zarkon):

you don't multiply the bases

OpenStudy (zarkon):

\[\Huge3^{-3}\times3^{4}\times3^{2}=3^{-3+4+2}=3^{3}\]

OpenStudy (zarkon):

\[3^{-3}\times3^{4}\times3^{2}=\frac{1}{3}\frac{1}{3}\frac{1}{3}3\cdot3\cdot3\cdot3\cdot3\cdot3\]

OpenStudy (zarkon):

\[=\frac{1}{\cancel{3}}\frac{1}{\cancel{3}}\frac{1}{\cancel{3}}\cancel{3}\cdot\cancel{3}\cdot\cancel{3}\cdot3\cdot3\cdot3\] \[=3^3\]

563blackghost (563blackghost):

O,o Excuse me, I must now go back and study multiplying exponents with common bases....

OpenStudy (study_buddy99):

according to mathway @563blackghost you're right

563blackghost (563blackghost):

@Zarkon is correct.... His simplification \(\Large{3^{3}=27}\) so Zarkon is correct...

OpenStudy (study_buddy99):

but... 3^3 is equal to 27 but 27 is in simplist form

OpenStudy (study_buddy99):

so wouldn't the answer be 27

563blackghost (563blackghost):

Yes it would be \(\LARGE{27}\) but when I solved it I had put up \(\LARGE{27^{3}}\) which is incorrect...

OpenStudy (study_buddy99):

oh!

563blackghost (563blackghost):

yup

OpenStudy (study_buddy99):

but what if they were to have different bases?